# GRADEO: Towards Human-Like Evaluation for Text-to-Video Generation via Multi-Step Reasoning

### v1

@@ 4.1 Experimental Setup
To thoroughly assess the correlation between model predictions and human evaluations, we perform experimental comparisons on both the test set and the constructed pairwise dataset. We extract 340 samples from the collected dataset GRADEO-Instruct to serve as the test set. On this dataset, we employ Spearman’s rank correlation coefficient (SROCC), Pearson’s linear correlation coefficient (PLCC), and Mean Absolute Error ($\bar{\Delta}$) as test metrics, calculating each dimension separately. SROCC evaluates rank correlation, while PLCC assesses the linear relationship, with higher values indicating stronger alignment with human evaluations, and a lower $\bar{\Delta}$ signifying a closer proximity between predicted and actual scores. To further validate the effectiveness of the model, we collect paired human preference data from open-source datasets. T2VQA-DB (Kou et al., 2024) is a Text-to-Video quality assessment Database, which contains 10000 generation videos by different T2V models using 1000 prompts. GenAI-Bench-Video (Li et al., 2024) collected 1.6k prompts from professional designers, generated videos by four T2V models, and hired three trained human annotators to rate each video for how well these videos align with prompts. TVGE (Wu et al., 2024) collected 2.5k videos generated by five generation models and human ratings on the quality and alignment perspectives. For T2VQA-DB, GenAI-Bench-Video, TVGE-Quality and TVGE-Alignment, we sample 200 prompts randomly, and for each prompt, select two videos, with a total of 1600 videos for pairwise comparisons. For each pair, one video has a higher score than the other. The accuracy of preference judgments, i.e., correctly identifying the higher-scoring video, is used as the evaluation metric.
@@
On the test dataset, we primarily selecte current state-of-the-art image and video multimodal large language models. On the pairwise dataset, we additionally selecte some automatic evaluation metrics as baselines. 1) Image MLLMs: LLaVa-1.5 (Liu et al., 2023), LLaVa-1.6 (Liu et al., 2024a), GLM-4v (GLM et al., 2024), GPT-4-Turbo, GPT-4o (Achiam et al., 2023). 2) Video MLLMs: InternVL-2.5 (Chen et al., 2024b), LLaVA-NeXT-Video (Zhang et al., 2024b), Gemini-1.5-Flash, Gemini-1.5-Pro (Team et al., 2023), VideoScore-1.1 (He et al., 2024). 3)Automatic Metrics: PIQE (N et al., 2015), BRISQUE (Mittal et al., 2012), CLIP-Score (Hessel et al., 2021), ImageReward (Xu et al., 2023), HPS-2.1 (Wu et al., 2023b). For MLLMs, with the exception of VideoScore, the input text for the other models are consistent with those of GRADE, comprising task-specific instructions and dimensional evaluation criteria. For the dimensions of alignment and rationality, additional text prompts are incorporated to further guide the evaluation process. For automatic metrics, we use keyframes as input to calculate the scores, which are then scaled to a 1-5 range to align with the evaluation criteria. For VideoScore, we map visual quality, temporal consistency, text-to-video alignment, and factual consistency to the four dimensions of quality, consistency, alignment, and rationality, respectively. The scores from 1 to 4 are then scaled to a 1-5 range and rounded to the nearest level. More details are shown in Appendix C.
@@
We adopt Qwen2-VL-7B (Wang et al., 2024a) as our base model, which has proven its perception and understanding capabilities for multimodal data, particularly videos, across various open-source datasets and downstream tasks. To enhance the model’s ability to assess generated videos in a manner similar to human evaluations, while preserving its original instruction-following and reasoning capabilities, we apply the LoRA fine-tuning method (Hu et al., 2021). The learning rate is set to $1\times 10^{-5}$, and the model is trained for 10 epochs on 4 RTX 3090 (24G) GPUs.
@@
PIQE (N et al., 2015) 0.300 0.220 0.305 0.185
BRISQUE (Mittal et al., 2012) 0.420 0.240 0.360 0.295
CLIP-Score (Hessel et al., 2021) 0.370 0.445 0.590 0.425
ImageReward-v1.0 (Xu et al., 2023) 0.490 0.640 0.730 0.605
HPS-v2.1 (Wu et al., 2023b) 0.475 0.730 0.660 0.625
@@
Gemini-1.5-Flash (Team et al., 2023) 0.470 0.720 0.620 0.460
VideoScore-v1.1 (He et al., 2024) 0.790 0.760 0.690 0.795
@@
Figure 4: Qualitative Results. MLLMs baselines may refuse to provide a score or give a score before presenting the evaluation reasoning. In contrast, our approach generates detailed reasoning steps prior to assigning a final score, resulting in higher consistency with human assessments.
@@ 4.2 Experimental Results
In addition to quantitative results, we also present qualitative results in Figure 4. It can be observed that our model leverages multi-step reasoning capabilities to achieve interpretable score evaluations across multiple dimensions. We have found that some MLLMs occasionally refuse to answer or provide very brief responses, which often leads to evaluation failures or errors. LLava-1.5-7b and Gemini-1.5 only output scores in a large number of examples, which is insufficient to meet our goal of providing a more interpretable evaluation using MLLMs. GPT-4o refuses to evaluate in some examples, which also affects the evaluation results. Models like Llava-NeXT-Video-7b typically respond with a score first, which leads to the evaluation rationale being generated based on that score, rather than utilizing reasoning capabilities to improve the evaluation results.
@@
4.3 Ablation Study
Fine-tuning w <Description> w <Overview> w CoT Human Correlation
$\bar{\rho_{srocc}}$($\uparrow$)/$\bar{\rho_{plcc}}$($\uparrow$)/$\bar{\Delta}$($\downarrow$)
$\times$ $\times$ $\times$ $\times$ 0.512/0.339/1.066
$\times$ $\times$ $\times$ 0.572/0.432/1.023
$\times$ $\times$ $\times$ 0.538/0.424/0.642
$\times$ $\times$ 0.588/0.462/0.610
0.692/0.647/0.428
Basic CNN Video Scoring Model 0.302/0.186/0.910
5 Benchmarking recent T2V models
5.1 Benchmark Setup
We adopt 8 open-source SOTA Text-to-Video models, including Hotshot-XL (Mullan et al., 2023), Lavie (Wang et al., 2024b), Latte (Ma et al., 2024), ModelScope (Wang et al., 2023), ZeroScope (Cerspense, 2024), VideoCrafter-1 (Chen et al., 2023), VideoCrafter-2 (Chen et al., 2024a), Open-Sora (Zheng et al., 2024). We evaluate models on GRADEO-Bench using GRADEO, and the scores obtained for each dimension are averaged and scaled to a percentage. Models setting and further details are provided in Appendix D.
@@ 6 Conclusion
This paper presents a novel evaluator and benchmark for evaluating video generation models. The development of GRADEO-Instruct and GRADEO fills a crucial gap in assessing AI-generated video content, establishing a foundation for future research in this field. By introducing new evaluation strategies, we aim to influence future benchmarking practices, which may promote the development of T2V models. However, the advancement of T2V models also brings potential negative impacts, including the risk of misuse to create realistic but misleading video content, such as deepfakes, which could contribute to misinformation and societal harm. Ethical considerations are central to this work, ensuring that the technology is developed and applied responsibly while adhering to rigorous standards. Our approach emphasizes that advancements in video generation models must be both beneficial and accountable to society.
@@ A.4 Prompts for Converting Instruction Tuning Data
We leverage GPT-4o to transform disorganized human feedback into logical and systematic reasoning CoT data. Three keyframes—initial, final, and a middle frame—are extracted as image inputs, while text prompts are illustrated in Figures 10, 11, and 12. Figure 10 represents the general prompt, while Figures 11 and 12 correspond to the Alignment and Rationality dimensions, respectively, as these dimensions require textual alignment and specialized transformations.
@@ C.1 Baselines Settings
C.2 More Qualitative Results
In this section, we present additional qualitative results to demonstrate how our method generates systematic reasoning steps to perform comprehension and commonsense reasoning, achieving better evaluation outcomes. In contrast, other SOTA MLLMs, not fine-tuned for video evaluation tasks, exhibit various limitations.(See Figure 13)
Figure 13: More qualitative results.
C.3 Ablation Details
@@ D.2 Qualitative Comparison on T2V models
T2V models were evaluated under standardized conditions with the following video generation parameters: Hotshot-XL at 672x384 resolution, 8 fps, and 1.0 second duration; Open-Sora at 424x240 resolution, 24 fps, and 4.25 seconds; VideoCrafter-2 and VideoCrafter-1 both at 512x320 resolution, 10 fps, and 1.6 seconds; ZeroScope at 576x320 resolution, 10 fps, and 2.4 seconds; ModelScope at 256x256 resolution, 8 fps, and 2.0 seconds; Lavie at 512x320 resolution, 8 fps, and 2.0 seconds; and Latte at 512x512 resolution, 10 fps, and 1.6 seconds. By generating videos using various T2V models on the Prompt Suite and scoring them with our approach, we present some qualitative comparisons, as shown in Figure 16, 17, 18, and 19.

### v2

@@ 4.1 Experimental Setup
To thoroughly assess the correlation between model predictions and human evaluations, we perform experimental comparisons on both the test set and the constructed pairwise dataset. We extract 340 samples from the collected dataset GRADEO-Instruct to serve as the test set. On this dataset, we employ Spearman’s rank correlation coefficient (SROCC), Pearson’s linear correlation coefficient (PLCC), and Mean Absolute Error ($\bar{\Delta}$) as test metrics, calculating each dimension separately. SROCC evaluates rank correlation, while PLCC assesses the linear relationship, with higher values indicating stronger alignment with human evaluations, and a lower $\bar{\Delta}$ signifying a closer proximity between predicted and actual scores. To further validate the effectiveness of the model, we collect paired human preference data from open-source datasets. T2VQA-DB (kou2024subjective) is a Text-to-Video quality assessment Database, which contains 10000 generation videos by different T2V models using 1000 prompts. GenAI-Bench-Video (li2024genai) collected 1.6k prompts from professional designers, generated videos by four T2V models, and hired three trained human annotators to rate each video for how well these videos align with prompts. TVGE (wu2024towards) collected 2.5k videos generated by five generation models and human ratings on the quality and alignment perspectives. For T2VQA-DB, GenAI-Bench-Video, TVGE-Quality and TVGE-Alignment, we sample 200 prompts randomly, and for each prompt, select two videos, with a total of 1600 videos for pairwise comparisons. For each pair, one video has a higher score than the other. The accuracy of preference judgments, i.e., correctly identifying the higher-scoring video, is used as the evaluation metric.
@@
On the test dataset, we primarily selecte current state-of-the-art image and video multimodal large language models. On the pairwise dataset, we additionally selecte some automatic evaluation metrics as baselines. 1) Image MLLMs: LLaVa-1.5 (llava), LLaVa-1.6 (liu2024llavanext), GLM-4v (glm2024chatglm), GPT-4-Turbo, GPT-4o (gpt4). 2) Video MLLMs: InternVL-2.5 (chen2024internvl), LLaVA-NeXT-Video (zhang2024llavanext-video), Qwen2-VL-7B (Qwen2VL), Gemini-1.5-Flash, Gemini-1.5-Pro (team2023gemini), VideoScore-1.1 (he2024videoscore). 3)Automatic Metrics: PIQE (piqe), BRISQUE (BRISQUE), CLIP-Score (hessel2021clipscore), ImageReward (xu2023imagereward), HPS-2.1 (wu2023human). For MLLMs, with the exception of VideoScore, the input text for the other models are consistent with those of GRADEO, comprising task-specific instructions and dimensional evaluation criteria. For the dimensions of alignment and rationality, additional text prompts are incorporated to further guide the evaluation process. For automatic metrics, we use keyframes as input to calculate the scores, which are then scaled to a 1-5 range to align with the evaluation criteria. For VideoScore, we map visual quality, temporal consistency, text-to-video alignment, and factual consistency to the four dimensions of quality, consistency, alignment, and rationality, respectively. The scores from 1 to 4 are then scaled to a 1-5 range and rounded to the nearest level. More details are shown in Appendix C.
@@
We adopt Qwen2-VL-7B (Qwen2VL) as our base model, which has proven its perception and understanding capabilities for multimodal data, particularly videos, across various open-source datasets and downstream tasks. To enhance the model’s ability to assess generated videos in a manner similar to human evaluations, while preserving its original instruction-following and reasoning capabilities, we apply the LoRA fine-tuning method (hu2021lora). The learning rate is set to $1\times 10^{-5}$, and the model is trained for 10 epochs on 4 RTX 3090 (24G) GPUs.
@@
PIQE (piqe) 0.300 0.220 0.305 0.185
BRISQUE (BRISQUE) 0.420 0.240 0.360 0.295
CLIP-Score (hessel2021clipscore) 0.370 0.445 0.590 0.425
ImageReward-v1.0 (xu2023imagereward) 0.490 0.640 0.730 0.605
HPS-v2.1 (wu2023human) 0.475 0.730 0.660 0.625
@@
Gemini-1.5-Flash (team2023gemini) 0.470 0.720 0.620 0.460
VideoScore-v1.1 (he2024videoscore) 0.790 0.760 0.690 0.795
@@
Figure 4: Qualitative Results for Creativity Dimension. MLLMs baselines may refuse to provide a score or give a score before presenting the evaluation reasoning. In contrast, our approach generates detailed reasoning steps prior to assigning a final score, resulting in higher consistency with human assessments.
@@ 4.2 Experimental Results
In addition to quantitative results, we also present qualitative results in Figure 4, 5. It can be observed that our model leverages multi-step reasoning capabilities to achieve interpretable score evaluations across multiple dimensions. We have found that some MLLMs occasionally refuse to answer or provide very brief responses, which often leads to evaluation failures or errors. LLava-1.5-7b and Gemini-1.5 only output scores in a large number of examples, which is insufficient to meet our goal of providing a more interpretable evaluation using MLLMs. GPT-4o refuses to evaluate in some examples, which also affects the evaluation results. Models like Llava-NeXT-Video-7b typically respond with a score first, which leads to the evaluation rationale being generated based on that score, rather than utilizing reasoning capabilities to improve the evaluation results. For video Rationality evaluation, MLLMs without instruction tuning exhibit a higher susceptibility to prompt-induced hallucinations, where generated responses are disproportionately influenced by the input prompts rather than grounded in the visual content.
4.3 Ablation Study
Fine-tuning w <Description> w <Overview> w CoT Human Correlation
[…]
$\times$ $\times$ $\times$ $\times$ 0.512/0.339/1.066
$\times$ $\times$ $\times$ 0.572/0.432/1.023
$\times$ $\times$ $\times$ 0.538/0.424/0.642
$\times$ $\times$ 0.588/0.462/0.610
0.692/0.647/0.428
Basic CNN Video Scoring Model 0.302/0.186/0.910
Figure 5: Qualitative Results for Rationality Dimension. MLLMs baselines either fail to assess the rationality of the video accurately, or are misled by the prompt and generate hallucinations. Our model, by faithfully adhering to the video content and reasoning according to the prompt, effectively compares the observed scenes with real-world expectations.
5 Benchmarking recent T2V models
5.1 Benchmark Setup
We adopt 8 open-source SOTA Text-to-Video models, including Hotshot-XL (hotshotxl), Lavie (wang2023lavie), Latte (ma2024latte), ModelScope (wang2023modelscope), ZeroScope (zeroscope), VideoCrafter-1 (chen2023videocrafter1), VideoCrafter-2 (chen2024videocrafter2), Open-Sora (opensora). We evaluate models on GRADEO-Bench using GRADEO, and the scores obtained for each dimension are averaged and scaled to a percentage. Further details are provided in Appendix D.
@@ 5.3 Insights
Models with strong instruction-following capabilities and extensive training data are more prone to generating unsafe content.
Through our evaluation of several advanced video generation models, we observe that models with high instruction-following fidelity tend to more strictly adhere to prompt specifications, even when such prompts imply or explicitly request unsafe, unethical, or harmful content. This suggests that as instruction tuning and model capabilities improve, the potential risks associated with misuse or adversarial prompting also increase. Notably, these models often lack robust safety constraints or alignment mechanisms to reject or sanitize unsafe inputs, making them susceptible to generating content that violates ethical or safety norms. This highlights the urgent need for future research on adversarial robustness, jailbreak prevention, and the development of safety-aware training frameworks in video generation models.
@@ 6 Conclusion
Acknowledgements
We would like to thank the anonymous reviewers for their insightful comments and suggestions, which helped improve the quality of this paper. This work was partly supported by the National Natural Science Foundation of China (Nos.62311530100&62171251) and the Special Foundations for the Development of Strategic Emerging Industries of Shenzhen (No.KJZD20231023094700001).
@@ Acknowledgements
This paper presents a novel evaluator and benchmark for evaluating video generation models. The development of GRADEO-Instruct and GRADEO fills a crucial gap in assessing AI-generated video content, establishing a foundation for future research in this field. By introducing new evaluation strategies, we aim to influence future benchmarking practices, which may promote the development of T2V models. Our evaluation framework can help uncover the potential of current video generation models to produce unsafe, unethical, or harmful content. However, the advancement of T2V models also brings potential negative impacts, including the risk of misuse to create realistic but misleading video content, such as deepfakes, which could contribute to misinformation and societal harm. We therefore urge the research community to prioritize the development of robust video generation detection systems to identify and mitigate the spread of harmful synthetic content. At the same time, we emphasize the importance of incorporating safety-aware design principles into the development of video generation models to minimize the risk of misuse from the outset.
@@ A.4 Prompts for Converting Instruction Tuning Data
We leverage GPT-4o to transform disorganized human feedback into logical and systematic reasoning CoT data. Three keyframes—initial, final, and a middle frame—are extracted as image inputs, while text prompts are illustrated in Figures 11, 12, and 13. Figure 11 represents the general prompt, while Figures 12 and 13 correspond to the Alignment and Rationality dimensions, respectively, as these dimensions require textual alignment and specialized transformations.
@@ C.1 Baselines Settings
C.2 Ablation Details
@@ D.2 Qualitative Comparison on T2V models
T2V models were evaluated under standardized conditions with the following video generation parameters: Hotshot-XL at 672x384 resolution, 8 fps, and 1.0 second duration; Open-Sora at 424x240 resolution, 24 fps, and 4.25 seconds; VideoCrafter-2 and VideoCrafter-1 both at 512x320 resolution, 10 fps, and 1.6 seconds; ZeroScope at 576x320 resolution, 10 fps, and 2.4 seconds; ModelScope at 256x256 resolution, 8 fps, and 2.0 seconds; Lavie at 512x320 resolution, 8 fps, and 2.0 seconds; and Latte at 512x512 resolution, 10 fps, and 1.6 seconds.
By generating videos using various T2V models on the Prompt Suite and scoring them with our approach, we present some qualitative comparisons, as shown in Figure 16, 17, 18, and 19.